# Multitask and Multimodal Neural Tuning for Large Models

### v1

@@ 4 MMUD Benchmark
We constructed our dataset using open-source datasets: RefCOCO Yu et al. (2016), RefCOCOg Mao et al. (2016), and RefClef Nagaraja et al. (2016); Mao et al. (2016). Since these datasets only provide referring segmentation masks, we augmented them with annotations for other tasks. The dataset generation pipeline is illustrated in Figure 3. Initially, we employed GPT-4v to generate image captions describing the contents. Subsequently, GPT-4v was used to generate complex questions along with answers based on the image contents. We then performed manual filtration to ensure that the generated contents were meaningful and suitable for multimodal understanding. We filtered out samples in cases where: 1) the generated contents were meaningless; 2) the length of the generated contents was either too long or too short; 3) the answers could be directly inferred from the questions without reference to the images. To enhance the capability of complex reasoning, we manually inserted the <OBJ-i> token after each object in the answer, where $i$ represents the $i$-th object in the image. For example, when generating the question Can you judge the relationship between the two people in the front based on what happened in the picture?" with the answer with answer In this image, one of the women in the front is cutting a pizza, and a little girl is next to her…, we manually inserted the <OBJ> token after the expression of each object: In this image, one of the women in the front <OBJ-1> is cutting a pizza <OBJ-2>, and a little girl <OBJ-3> is next to her…. Following the structure of the RefCOCO datasets, each sample image contains multiple meaningful objects with corresponding masks. Each <OBJ-i> token not only corresponds to an object but also to the corresponding segmentation mask, which could help models better capture the relationship between texts and images. The dataset will be made publicly available to advance research in this area.
@@ 5.2 Results on MMUD
Table 1 and Table 2 presents the results of our proposed neural tuning compared with previous methods. Some visualized cases are shown in Figure  4 (Appendix A). For vanilla referring segmentation and reasoning segmentation, we mainly compare our method with LVAT, LISA, and PixelLM. We employ the mean intersection over union (mIoU) and overall IoU as the metrics. LVAT aims to fuse the BERT features into vision backbones, achieving great results in vanilla segmentation. However, when it comes to complex reasoning scenarios, LVAT fails to converge while other LLM-based methods yield good results, demonstrating the powerful reasoning capabilities of large models. Compared with LISA and PixelLM, which employ pretrained LLaVA Liu et al. (2023) and focus on complex reasoning segmentation, neural tuning can achieve better results, demonstrating its effectiveness. Regarding inference speed, our tuned 7B model can complete the reasoning segmentation process within 80ms per text-image pair on an RTX 4090 GPU, while the 13B model takes 110ms on an RTX A6000 GPU. This demonstrates the efficiency of our proposed neural tuning. We believe this efficiency results from the linear sparse vector updating and task guidance. Furthermore, to further evaluate the effectiveness of our proposed method, we also present the performance on the original test set of RefCOCO, RefCOCO+, and RefCOCOg datasets  [^2] For fair comparision, we re-split the train, validation, and test set of MMUD according to the original datasets.. As shown in Table 1(b), our method achieves state-of-the-art performance on the public datasets but with lower computational burden, revealing the effectiveness and efficiency of our proposed multitask neural tuning for segmentation tasks.
@@
For image captioning, we compare our method with BLIP-2 Li et al. (2023a), ExpansionNetV2 Hu et al. (2022) and mPLUG Li et al. (2022). BLIP-2 proposed a Q-Former for multimodal interaction whlie mPLUG learn the relationship between modalities by cross-modality skip-connection. We employ the BLEU-4 Papineni et al. (2002), CIDEr Vedantam et al. (2015), and METEOR Banerjee and Lavie (2005) metrics for evaluation. The results are shown in Table 2. As we can draw from the results, we can reach the competitive performance compared with previous state-of-the-art approaches. On text-to-image synthesis tasks, quantitative metrics are shown in Table 2. We employ FID score Heusel et al. (2017), KID Score Bińkowski et al. (2018), and inception score (IS Salimans et al. (2016)) as the metrics. We compare our method with previous methods, including Parti Yu et al. (2022), GLIGEN Li et al. (2023b), and U-ViT-S/2 Bao et al. (2023). Compared with other methods, one of the significant advantage of our method is that neural tuning is designed for multitask tuning instead of a certain task.
@@ 5.3 Ablation Study
Modules in neural tuning: The quantitative results are presented in Table 3. We observed that removing sparse vector updating (Equation 2), which renders the tuning layers of the sparse task network independent of each other, leads to a drop in performance. This highlights the significance of interlinking different layers of tuning networks, a factor overlooked by prior tuning methods such as LoRA Hu et al. (2021) and (IA)3 Liu et al. (2022). Additionally, performance degrades when the SDR strategy is eliminated, indicating that all neurons are activated for all tasks. Finally, we conducted an ablation study on multitask learning. The results show that removing the multitask learning strategy led to a relative decrease in performance. Notably, the vanilla referring segmentation and image captioning tasks appear to contribute more to complex reasoning than the image generation tasks, which corresponds to the visualized results in Figure 2(b).
@@
The selection of tuning modules: In Section 3.1, we integrated task guidance into the LLMs for query and value, following established methods Hu et al. (2021). However, this approach is not the sole option. We conducted corresponding ablation studies, with the quantitative results presented in Table 2. The results indicate that integrating task guidance into keys and values tends to result in poorer performance. While performance improves when task guidance is integrated into all query, key, and value modules, the improvement is only marginal compared to when it is integrated solely into query and value. Therefore, to strike a balance between performance and complexity, we opted to perform tuning only on query and value modules (Equation 5).
@@
Choice of image generation decoders: Apart from VQGAN, there are alternative options for image decoding. We also utilized a pre-trained DALL-E2 decoder [^3] http://github.com/hidden/for/review for generating images from the hidden embeddings. A comparison is presented in Figure A. The quality and style of the generated images are significantly influenced by the pre-trained decoders. This bias is a result of the pretraining data used for the decoders. Hence, in theory, neural tuning can employ any pre-trained image synthesis decoder. Unfortunately, due to hardware limitations, we were unable to test other pre-trained models for image generation.
6 Conclusion
In this paper, we introduce a novel tuning strategy called neural tuning. Neural tuning unifies tasks into an all-in-token approach, enhancing scalability for additional modalities or tasks. Furthermore, we emulate human thought processes using sparse distributed representation by activating specific neurons for different tasks. We evaluate our method across four tasks, including reasoning segmentation and text-to-image synthesis, demonstrating competitive performance compared to current state-of-the-art methods. To further research in this field, we introduce the MMUD dataset, which includes various annotations for different tasks.
7 Limitations
One of the biggest limitation is that our research currently does not incorporate acoustic tasks into neural tuning and MMUD. Ideally, under optimal conditions, neural tuning could accommodate all types of multimodal tasks by appropriately formulating them into tokens. In the future, we aim to conduct more studies on multitask learning for large multimodal models, exploring new technologies, models, and benchmarks. Additionally, similar to other large-scale models, this work carries the potential risk of misuse. To mitigate this issue, we will implement an application-based access system for the release of data and models. This approach aims to ensure their use remains within reasonable and ethical boundaries.
Appendix A Case Visualization
The visualization of some cases from our model are shown in Figure A. The tasks included in these cases include: vanilla referring segmentation, reasoning segmentation, image captioning, and image generation. For image generation, we employ two pretrained decoders: VQGAN and DALL-E2.
@@ Appendix A Case Visualization
Appendix B Ablation of $\beta$ in Sparse Vector Updating
The detailed ablations regarding $\beta$ in sparse vector updating are presented in Table 4. In these experiments, we consistently use the reasoning segmentation task and employ the corresponding oIoU for evaluation. For instance, two tasks involved refer to reasoning segmentation and vanilla segmentation, while only one task involved refers to reasoning segmentation only. From the results, we observe that setting $\beta$ to 0.4 yields the best performance when all four tasks are involved. For scenarios involving three tasks, the optimal performance is achieved with $\beta$ set to 0.5 or 0.6. Finally, when there are only one or two tasks, it is more beneficial to activate a higher proportion of neurons, with $\beta$ set to 0.9 or 1.0.
@@ Appendix B Ablation of $\beta$ in Sparse Vector Updating
Appendix C Ablation of Activation Patterns for Sparse Vector $Z_{k}^{u}$
Although we have conducted experiments on the activation rate $\beta$ of the sparse vector $Z_{k}^{u}$ (Appendix B), the activation pattern itself is still worth exploring. In designing our activation strategy, we aimed to activate different neurons according to different tasks and allow the number of activations to fluctuate slightly, mirroring the human thinking process. The simplest and most intuitive approach is to predefine an activation rate and allow it to fluctuate according to a Gaussian distribution, as illustrated in Section 3.1. However, other patterns can also adhere to this activation principle. Below are additional activation strategies explored in our experiments (assuming we need to activate $\beta\%$ of neurons and the fluctuation rate is $f\%$, where $f<\beta$):

### v2

@@ 4 MMUD Benchmark
We constructed our dataset using open-source datasets: RefCOCO [27], RefCOCOg [28], and RefClef [29, 28]. Since these datasets only provide referring segmentation masks, we augmented them with annotations for other tasks. The dataset generation pipeline is illustrated in Figure 3. Initially, we employed GPT-4v to generate image captions describing the contents. Subsequently, GPT-4v was used to generate complex questions along with answers based on the image contents. We then performed manual filtration to ensure that the generated contents were meaningful and suitable for multimodal understanding. We filtered out samples in cases where: 1) the generated contents were meaningless; 2) the length of the generated contents was either too long or too short; 3) the answers could be directly inferred from the questions without reference to the images. To enhance the capability of complex reasoning, we manually inserted the $<$OBJ-i$>$ token after each object in the answer, where $i$ represents the $i$-th object in the image. For example, when generating the question Can you judge the relationship between the two people in the front based on what happened in the picture?” with the answer with answer In this image, one of the women in the front is cutting a pizza, and a little girl is next to her…, we manually inserted the $<$OBJ$>$ token after the expression of each object: In this image, one of the women in the front $<$OBJ-1$>$ is cutting a pizza $<$OBJ-2$>$, and a little girl $<$OBJ-3$>$ is next to her…. Following the structure of the RefCOCO datasets, each sample image contains multiple meaningful objects with corresponding masks. Each $<$OBJ-i$>$ token not only corresponds to an object but also to the corresponding segmentation mask, which could help models better capture the relationship between texts and images. The dataset will be made publicly available to advance research in this area.
@@ 5.2 Results on MMUD
Table 1 and Table 2 presents the results of our proposed neural tuning compared with previous methods. For vanilla referring segmentation and reasoning segmentation, we mainly compare our method with LVAT, LISA, and PixelLM. We employ the mean intersection over union (mIoU) and overall IoU as the metrics. LVAT aims to fuse the BERT features into vision backbones, achieving great results in vanilla segmentation. However, when it comes to complex reasoning scenarios, LVAT fails to converge while other LLM-based methods yield good results, demonstrating the powerful reasoning capabilities of large models. Compared with LISA and PixelLM, which employ pretrained LLaVA [12] and focus on complex reasoning segmentation, neural tuning can achieve better results, demonstrating its effectiveness. Regarding inference speed, our tuned 7B model can complete the reasoning segmentation process within 80ms per text-image pair on an RTX 4090 GPU, while the 13B model takes 110ms on an RTX A6000 GPU. This demonstrates the efficiency of our proposed neural tuning. We believe this efficiency results from the linear sparse vector updating and task guidance.
Furthermore, to further evaluate the effectiveness of our proposed method, we also present the performance on the original test set of RefCOCO, RefCOCO+, and RefCOCOg datasets  [^2] For fair comparision, we re-split the train, validation, and test set of MMUD according to the original datasets.. As shown in Table 1(b), our method achieves state-of-the-art performance on the public datasets but with lower computational burden, revealing the effectiveness and efficiency of our proposed multitask neural tuning for segmentation tasks.
@@
For image captioning, we compare our method with BLIP-2 [10], ExpansionNetV2 [30] and mPLUG [31]. BLIP-2 proposed a Q-Former for multimodal interaction whlie mPLUG learn the relationship between modalities by cross-modality skip-connection. We employ the BLEU-4 [32], CIDEr [33], and METEOR [34] metrics for evaluation. The results are shown in Table 2. As we can draw from the results, we can reach the competitive performance compared with previous state-of-the-art approaches.
On text-to-image synthesis tasks, quantitative metrics are shown in Table 2. We employ FID score [35], KID Score [36], and inception score (IS [37]) as the metrics. We compare our method with previous methods, including Parti [38], GLIGEN [39], and U-ViT-S/2 [40]. Compared with other methods, one of the significant advantage of our method is that neural tuning is designed for multitask tuning instead of a certain task.
5.3 Case Visualization
The visualizations of several cases generated by our model are shown in Figure 4. These cases encompass a variety of tasks, including vanilla referring segmentation, reasoning segmentation, image captioning, and image generation. For image generation, we utilize two pretrained decoders: VQGAN and DALL-E2. In the case of reasoning segmentation, the model not only accurately segments the corresponding objects but also provides detailed reasoning or explanations for the results. These examples highlight the model’s ability to seamlessly handle diverse multimodal tasks, demonstrating both its precision in segmentation and its creative capabilities in image generation.
Figure 5: Visualization of Attention Scores for a Reasoning Segmentation Case. For better clarity, we omit the question and answer tokens, as well as some image patches. Darker regions indicate higher attention weights. The results demonstrate that the model effectively builds connections between image details, text, and the ¡OBJ¿ task token.
5.4 Effectiveness of Self-Attention in All-in-token Paradigm
In our approach, we utilize self-attention to compute the attention scores between the various elements of the input, including both image patches and textual tokens. This differs from traditional methods that rely on cross-attention, where separate attention mechanisms are used for different modalities. The self-attention mechanism, by contrast, enables a more unified interaction between modalities, allowing the model to capture both intra-modal and inter-modal relationships simultaneously.
As shown in Figure 5, we visualize the attention scores between image patches and corresponding textual tokens for reasoning segmentation with relatively shorter sentences. The figure demonstrates that object-related words in the text (such as ”cat,” ”car,” or ”tree”) receive significantly higher attention weights from their corresponding image patches. This indicates that the self-attention mechanism effectively focuses on the relevant visual regions when processing the corresponding linguistic cues, facilitating a more coherent integration of multimodal information. The ability to compute these attention scores in a unified manner not only streamlines the model’s architecture but also enhances its capacity to perform tasks requiring intricate cross-modal reasoning, such as referring segmentation and text-to-image synthesis.
@@ 5.5 Ablation Study
Modules in neural tuning: The quantitative results are presented in Table 3. We observed that removing sparse vector updating (Equation 2), which renders the tuning layers of the sparse task network independent of each other, leads to a drop in performance. This highlights the significance of interlinking different layers of tuning networks, a factor overlooked by prior tuning methods such as LoRA [41] and (IA)3 [42]. Additionally, performance degrades when the SDR strategy is eliminated, indicating that all neurons are activated for all tasks. Finally, we conducted an ablation study on multitask learning. The results show that removing the multitask learning strategy led to a relative decrease in performance. Notably, the vanilla referring segmentation and image captioning tasks appear to contribute more to complex reasoning than the image generation tasks, which corresponds to the visualized results in Figure 2(b).
@@
The selection of tuning modules: In Section 3.2, we integrated task guidance into the LLMs for query and value, following established methods [41]. However, this approach is not the sole option. We conducted corresponding ablation studies, with the quantitative results presented in Table 2. The results indicate that integrating task guidance into keys and values tends to result in poorer performance. While performance improves when task guidance is integrated into all query, key, and value modules, the improvement is only marginal compared to when it is integrated solely into query and value. Therefore, to strike a balance between performance and complexity, we opted to perform tuning only on query and value modules (Equation 5).
@@
Choice of image generation decoders: Apart from VQGAN, there are alternative options for image decoding. We also utilized a pre-trained DALL-E2 decoder [^3] http://github.com/hidden/for/review for generating images from the hidden embeddings. A comparison is presented in Figure 4 (Section  5.3). The quality and style of the generated images are significantly influenced by the pre-trained decoders. This bias is a result of the pretraining data used for the decoders. Hence, in theory, neural tuning can employ any pre-trained image synthesis decoder. Unfortunately, due to hardware limitations, we were unable to test other pre-trained models for image generation.
@@
Ablation of $\beta$ in Sparse Vector Updating The detailed ablations regarding $\beta$ in sparse vector updating are presented in Table 4. In these experiments, we consistently use the reasoning segmentation task and employ the corresponding oIoU for evaluation. For instance, two tasks involved refer to reasoning segmentation and vanilla segmentation, while only one task involved refers to reasoning segmentation only. From the results, we observe that setting $\beta$ to 0.4 yields the best performance when all four tasks are involved. For scenarios involving three tasks, the optimal performance is achieved with $\beta$ set to 0.5 or 0.6. Finally, when there are only one or two tasks, it is more beneficial to activate a higher proportion of neurons, with $\beta$ set to 0.9 or 1.0.
@@
Ablation of Activation Patterns for Sparse Vector $Z_{k}^{u}$ Although we have conducted experiments on the activation rate $\beta$ of the sparse vector $Z_{k}^{u}$ (Appendix5.5), the activation pattern itself is still worth exploring. In designing our activation strategy, we aimed to activate different neurons according to different tasks and allow the number of activations to fluctuate slightly, mirroring the human thinking process. The simplest and most intuitive approach is to predefine an activation rate and allow it to fluctuate according to a Gaussian distribution, as illustrated in Section 3.2. However, other patterns can also adhere to this activation principle. Below are additional activation strategies explored in our experiments (assuming we need to activate $\beta\%$ of neurons and the fluctuation rate is $f\%$, where $f<\beta$):
  • 1.
@@
6 Conclusion
In this paper, we present a novel unified multimodal multitask learning framework with a new tuning strategy called neural tuning. Under this framework, we unify tasks using an all-in-token approach, which enhances scalability by facilitating the integration of additional modalities or tasks. Additionally, we emulate human cognitive processes through sparse distributed representation, activating specific neurons for different tasks. We evaluate our method across four tasks, including reasoning segmentation and text-to-image synthesis, and demonstrate competitive performance compared to current state-of-the-art methods. To support further research in this area, we introduce the MMUD dataset, which provides a diverse set of annotations for various tasks. Moving forward, we plan to explore additional modalities, such as acoustics, and expand our research into fields like medical applications.
Appendix A Example Appendix Section
Appendix text.